OpenStudy (anonymous):

prove that f(x) = x^2 and g(x) = 1000x^2 grow at the same rate.

OpenStudy (anonymous):

@freckles @ganeshie8 @Hero @pooja195 any clues ?

ganeshie8 (ganeshie8):

If \(\lim\limits_{x\to \infty}\dfrac{f(x)}{g(x)} \) is a finite nonzero number, then we say that \(f(x)\) and \(g(x)\) grow at same rate.

OpenStudy (anonymous):

okay that helps loads :D

OpenStudy (anonymous):

okay im getting .001

OpenStudy (anonymous):

2/2000

ganeshie8 (ganeshie8):

which is finite and not 0, so both function grow at the same rate

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

yw